Witten Entertainment is considering buying a machine that costs $560,000. The machine will be depreciated over four years by the straight-line method and will be worthless at that time. The company can lease the machine with year-end payments of $173,000. The company can issue bonds at an interest rate of 7 percent. The corporate tax rate is 25 percent.

What is the NAL of the lease?
Answer #1

NAL $ -20,818.66. So, there is no advantage of leasing.

Step-1:Calculation of Present value of cash flow under leasing option
Year After Tax Lease Rental Lost depreciation tax shield Total Cash Flow Discount factor Present Value
a b c d=b+c e=1.0525^-a f=d*e
1 1,29,750.00 35,000.00 1,64,750.00 0.9501 1,56,532.07
2 1,29,750.00 35,000.00 1,64,750.00 0.9027 1,48,724.05
3 1,29,750.00 35,000.00 1,64,750.00 0.8577 1,41,305.51
4 1,29,750.00 35,000.00 1,64,750.00 0.8149 1,34,257.02
Total 5,80,818.66
Working:
After Tax lease rental = Lease rental * (1-Tax Rate)
= 173000*(1-0.25)
= 1,29,750
Straight Line depreciation = (Cost - Salvage Value)/Useful life
= (560000-0)/4
= 1,40,000
Depreciation Tax Shield = Depreciation Amount* Tax Rate
= 140000*25%
= 35000
After -Tax Discount rate = Interest rate*(1- Tax Rate)
= 7%*(1-0.25)
= 0.0525
Step-2:Calculation of net advantage of leasing
Cost of asset 5,60,000.00
Less Present value of cash flow under lease option 5,80,818.66
Net Advantage of Leasing -20,818.66